Technical Staff Blog

Last update on .

All tstaff managed linux desktop will be rebooted on Wednesday, August 31st, at 7am.

Tstaff is pushing out a new nvidia driver for all managed linux desktop running X in order support new nvidia graphics hardware.  Any managed linux desktop with the following nvidia cards listed below will get the new 367.44 driver.  A ...

Last update on .

whoa-1444580.jpg At about 1am this morning, disk hardware providing a backing store for our VMWare machines went offline. This caused nearly every one of our production servers to go offline as well as all our hosting class machines provisioned for users and research groups. CIS is investigating the issue and we will post updates as we ...

Last update on .

power-outage.jpg We experienced some sort of power blip on Saturday, likely the result of the storms that rolled through. This took out a number of grid machines. The majority of the machines are back and operational again, but obviously any jobs running on the machines have been killed. There are still about two dozen machines we ...

Last update on .

Unfortunately, there is an issue with the Grid Master process which has rendered qstat, qsub, etc... unusable.  We're working on diagnosing the problem, now.  But, our Grid Engine guru is out of town, at the moment.

Any jobs that were submitted before the outage are still running and will continue to run, as expected ...

Last update on .

CIS Networking Group announced CS edge switches maintenance schedule on August 2nd, 6am - 8am.  See below.   

During this maintenance windows, all desktop machine network connection will be affected.  This includes most of ssh forwarding hosts except a few virtualize hosts like unison, oolong, kui, emma, eternity, earl and bushtea. A few servers will also be ...

Last update on .

We had a quick 2 minute GPFS departmental cluster nfs blip at 8:27am when one of the cluster nfs server, nerds, experienced a kernel panic.  The failover of the nfs ip address to its partner was success.  We are debugging the issue with nerds and will bring it back into the GPFS departmental nfs ...

Last update on .

  •     New VPN configuration files are now available at software.cs.brown.edu.
  •     Old VPN configuration files will expire on July 10th.
  •     A short VPN outage is scheduled for 12pm on June 24th.

openvpn.png We were unable to renew the VPN server certificate on schedule, and we would like to apologize for any inconvenience that this may ...

Last update on .

CIS Networking group is planning to make changes to the 10G switches, cit-cs-r14-sw-1 and cit-cs-r14-sw-2.  The switch config change requires a reboot.  They will be changing and rebooting the cit-cs-r14-sw-1 on Tuesday, June 21st at 5:30am and cit-cs-r14-sw-2 on Thursday, June 23rd at 6:00am.  The GPFS NSD nodes are connected to these switches ...

Last update on .

  • New VPN configuration files are now available at software.cs.brown.edu.
  • Old VPN configuration files will expire on July 10th.
  • A short VPN outage is scheduled for 12pm on June 22nd.

 Tstaff will be renewing the Certificates used to provide the SSL connection for our VPN service.  As such, the new .ovpn configuration files ...

Yearly archives